OpenStudy (anonymous):

A biologist is comparing the growth of a population of flies per week to the number of flies a bullfrog will consume per week. She has devised an equation to solve for which day (x) the bullfrog would be able to eat the entire population. The equation is 2x = 3x - 1. However, she has observed that the bullfrog cannot eat more than seven flies in one week. Explain to the biologist how she can solve this on a graph using a system of equations. Identify any possible constraints to the situation.

OpenStudy (compassionate):

Is there any additional info? Why two x's? Should be an x and y

OpenStudy (anonymous):

That's the whole question and that's why I'm confused :/

OpenStudy (compassionate):

Well, let me do some sloppy maths here. 2x = 3x - 1 -x = -1 x = 1 So then plug x in 2(1) = 3(1) - 1 2 = 2 So it checks out okay. But this makes no sense, because he will consume the entire population on day 1. Something is very wrong.

OpenStudy (anonymous):

Well the bullfrog can't eat more than 7

OpenStudy (compassionate):

Are you sure the equation isn't 2^x = 3x - 1? Check it again.

OpenStudy (anonymous):

you're correct.

OpenStudy (anonymous):

it's 2^x = 3x - 1

OpenStudy (anonymous):

@Compassionate I still can't manage to understand this :/

OpenStudy (compassionate):

Okay. Take 2^2 = 3x-1 and put it in two different equations. y = 2^x and y = 3x - 1 You intersect at 1 and 3 Do you get where I'm going with this?

OpenStudy (anonymous):

I get why we're splitting it into two different equations.

OpenStudy (anonymous):

I'm pretty confused to be honest. (bad at math)

OpenStudy (compassionate):

Okay, so you understand that we split it into two equations to simplify it. Right? Now simply graph each one.

OpenStudy (anonymous):

can you "Explain to the biologist how she can solve this on a graph using a system of equations." I'm understanding the graphing, I just don't know what they want me to tell them.

OpenStudy (compassionate):

We can solve x below y=2x and y=3x−x these = graphs give us the x(x) So the intersection happens at x = 1 and x = 3 She said the bullfrog cannot eat more seven a day So, it will only take him 1 day.

OpenStudy (anonymous):

Thanks :)